OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y?

OpenStudy (anonymous):

2x - 3y = - 10 when x = 4 substitute 4 in for x..... 2(4) - 3y = - 10 8 - 3y = - 10 (get y on one side) -3y = - 10 - 8 - 3y = - 18 y = 6